# Jina-ColBERT-v2: A General Purpose Multilingual Late Interaction Retriever

### v1

@@ 7.2 Query Augmentation Attention
An important feature of ColBERT’s implementation is its query augmentation mechanism. By padding queries with [MASK] tokens to a uniform length, ColBERT uses BERT’s masked language modeling ability to produce additional soft term embeddings which interact with document token embeddings during MaxSim scoring. However, prior ColBERT models do not modify the attention mask to allow query tokens to attend to the mask tokens, which some hypothesize might harm generalization by making this augmentation feature too integral to the embedding process. Our controlled triplet training experiment in Table 2, however, demonstrates a positive effect across a variety of tasks, with particular benefit to non-English tasks in MIRACL.
Table 2: nDCG@10 scores for Score Normalization and [MASK] attention experiments on a retrieval-oriented subset of BEIR and MIRACL tasks. *Results for score normalization are performed on the BM25-sampled versions of all datasets presented except TREC-COVID (tc), while [MASK] attention experiments report full-fidelity scores.
7.3 Supervision Loss
Following ColBERTv2, we fine-tune Jina-ColBERT-v2’s pair-trained checkpoint on samples with hard negatives using a KL divergence loss function to distill soft labels from the teacher model. For the teacher model, we use jina-reranker-v2-base-multilingual [^3] https://huggingface.co/jinaai/jina-reranker-v2-base-multilingual, a highly capable multilingual cross encoder.
7.4 Score Normalization
Recently, clavie_jacolbertv25_2024 applied min-max normalization to both the student and teacher scores before computing the KL loss. This adjustment brings the score distributions of the ColBERT model and its CE teacher into closer alignment, as the original score distribution for ColBERT theoretically ranges from zero to the number of query tokens, and is model-dependent for the teacher CE. Our experiment presented in Table 2, however, shows this method to have inconclusive benefit to nDCG@10 on the BEIR and MIRACL datasets when applied to our model. We consider this result to be understandable given clavie_jacolbertv25_2024’s very small observed effect.
8 Results
In this section we compare Jina-ColBERT-v2 to a variety of models across the BEIR, LoTTE, MIRACL, and mMARCO benchmarks to situate it’s performance in the landscape of other monolingual and multilingual retrievers.
Table 4 shows Jina-ColBERT-v2’s strong English performance compared to ColBERTv2, averaging a 6.6% improvement. [TODO: if we include answerai-small: while still trailing answerai-colbert-small by 4.8%.] Notably, however, we perform well below ColBERTv2 on ArguAna, which we might attribute to either its unusual task: counterargument retrieval being at odds with our retrieval-heavy triplet training data distribution, or as an indication of the limitation of our stronger augmentation attention, discussed in Section 7.2, when applied to much longer (300 token) queries. Similarly for LoTTE, we see in table LABEL:tab:lotte_final a 6.1% average improvement over ColBERTv2 the five search domains.
Table LABEL:tab:miracl_final compares Jina-ColBERT-v2 to Bm25, mDPR, and BGE-M3. While we handily outperform BM25 and zhang2023miracl-reported zero-shot and in-langauge-finetuned mDPRs as expected, our model is slightly outperformed (-0.66%) by the better-finetuned mDPR from (zhang_2023_towards), and heavily by BGE-M3 (-9.31%). These comparisons, however, must acknowledge that BGE-M3’s token representations are 8 times the size of Jina-ColBERT-v2’s ($128\to 1024$), and that the mDPR model is only finetuned for one language, rather than many like ours which may suffer to some extent from the curse of multilinguality.
ar de dt es en fr hi id it ja pt ru vi zh
ColBERT-XM 19.5 27.0 27.5 28.5 37.2 26.9 23.8 26.3 26.5 24.1 27.6 25.1 22.6 24.6
Jina-ColBERT-v2 27.2 33.1 33.0 34.1 39.6 33.5 30.9 31.9 33.7 27.6 33.7 29.8 28.7 30.2
Table 3: Comparison of mRR@10 scores between ColBERT-XM and Jina-ColBERT-v2 models on the mMARCO dev set across various languages.
Table 4: Comparison of nDCG@10 scores between BM25, ColBERTv2, answer-colbert-small and Jina-ColBERT-v1 and Jina-ColBERT-v2 on the BEIR test set.
@@ 9 Conclusion
This work presents Jina-ColBERT-v2, a capable multilingual ColBERT model that is the result of improvements to its architecture and training process. We implement modifications to the model architecture like flash attention and MRL heads that yield efficiency gains with effectively no downside, and subsequently train it on a heterogenous mix of data of varying tasks, languages, and supervision structures in order to bolster its performance as a general purpose retriever.
@@
Appendix A Appendix
Figure 2: Data distribution after sampling for non-English languages. [TODO: to appendix?]
Figure 3: Data distribution after sampling for triplet datasets. [TODO: to appendix?]

### v2

@@ 8.4 Query Augmentation Attention
BEIR MIRACL
tc hp nq ms de es fr ja zh
with query augmentation 80.2 71.5 58.8 44.3 45.6 49.8 44.8 58.8 52.9
without query augmentation. 77.2 70.4 54.6 37.6 33.3 40.3 35.9 54.9 34.4
Table 7: nDCG@10 scores for the query augmentation attention experiments on a retrieval-oriented subset of BEIR and MIRACL tasks. Results report full-fidelity scores.
An important feature of ColBERT’s implementation is its query augmentation mechanism. By padding queries with [MASK] tokens to a uniform length, ColBERT uses BERT’s masked language modeling ability to produce additional soft term embeddings which interact with document token embeddings during MaxSim scoring. However, prior ColBERT models do not modify the attention mask to allow query tokens to attend to the mask tokens, which some hypothesize might harm generalization by making this augmentation feature too integral to the embedding process. Our controlled triplet training experiment in Table 7, however, demonstrates a positive effect across a variety of tasks, with particular benefit to non-English tasks in MIRACL. We therefore allow this attention in our training and inference.
@@ 9 Conclusion
This work presents Jina-ColBERT-v2, a capable multilingual ColBERT model that is the result of improvements to its architecture and training process. We implement modifications to the model architecture like RoPE, flash attention and MRL heads that yield efficiency gains with effectively no downside, and subsequently train it on a heterogeneous mix of data of varying tasks, languages, and supervision structures in order to bolster its performance as a general purpose retriever.
@@
10 Acknowledgement
We thank Qi Liu and Jiaxin Mao from Renmin University of China for the contributions to Jina-ColBERT-v1 and offer the insights about MRL over MRL-E for ColBERT models.